OpenStudy (anonymous):

Please help! If x varies inversely with y and x = 6 when y = 10, find y when x = 8. A. y= 4/15 B. y= 4.8 C. y= 7.5 D. y= 40/3

OpenStudy (goformit100):

B. y= 4.8

OpenStudy (anonymous):

@goformit100 how did you get that?

OpenStudy (goformit100):

Can I please Show the working ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

y = k/x 10 = k/6 60 = k y = 60/8 y = 7.5 K is a constant here.

OpenStudy (goformit100):

Mr. @iambatman did correct.

OpenStudy (anonymous):

alright. thanksss